OpenStudy (anonymous):

Here's a limit question lim as x approaches 0 xsin(3/x)+34 This looks like a squeeze theorem problem, is that right?

OpenStudy (aravindg):

The first term becomes 0. I don't exactly remember the theorem name but x times bounded function when x moves to zero gives zero as an answer.

OpenStudy (anonymous):

Yeah the limit as x approaches 0 of xsin(1/x) (or any number over x inside the sin) =0. So does that not mean that the limit as x approaches 0 of xsin(3/x)+34 be equal to 34? Because essentially you just do (xsin(3/x)=0)+34=34?

OpenStudy (anonymous):

Yes I'm right, it's 34.